A body of 85kg moving with velocity of 9ms-¹ hits a stationary body of mass 60kg. If the two bodirs stick together after collision, calculate the magnitude of the velocity with whicj they move.

1
Expert's answer
2022-02-06T14:32:21-0500

Given:

"m_1=85\:\rm kg"

"v_1=9\:\rm m/s"

"m_2=60\:\rm kg"

"v_2=0\:\rm m/s"


The law of conservation of the momentum says

"m_1v_1+m_2v_2=(m_1+m_2)v_3"

Hence, the final velocity of system

"v_3=\frac{m_1v_1+m_2v_2}{m_1+m_2}"

"v_3=\frac{85*9+60*0}{85+60}=5.3\:\rm m/s"